# Double-coset zeta functions for groups acting on trees

### v1

@@ 6. The reciprocal of $\mathcal{Z}_{\Gamma,u\to u}(s)$
7. The behaviour at $s=-1$
The main goal of this section is to prove that the Euler–Poincaré identity for a unimodular t.d.l.c. group $G$ having a weakly locally $\infty$-transitive or (P)-closed action on a tree as prescribed by Corollary F. To achieve this result, we first give a formula of the relevant Euler–Poincaré characteristic in terms of local data of the action (cf. Proposition 7.3) and then use the splitting formulae of Section 6 to prove Theorem E and Corollary F. Finally, given an edge-weighted graph $(\Gamma,\omega)$, we exploit the machinery introduced in this paper to relate the behaviour of $\mathcal{Z}_{\Gamma,u\to u}(-1)$ with the behaviour at $1$ of a suitable weighted Ihara zeta function associated to $\Gamma$ (cf. Section 7.3).
@@ 7.1. The Euler–Poincaré characteristic
According to [8, §5], every unimodular t.d.l.c. group $G$ of type $\mathrm{FP}$ (with respect to the category of discrete left $\mathbb{Q}[G]$-modules) admits an Euler–Poincaré characteristic $\widetilde{\chi}_{G}$. For every compact open subgroup $K\leq G$, this invariant is a determined rational multiple of the Haar measure $\mu_{K}$ on $G$ normalised with respect to $K$, written $\widetilde{\chi}_{G}=\chi(G,\mu_{K})\cdot\mu_{K}$. By the uniqueness of the Haar measure on $G$ up to a positive real rescaling, for every Haar measure $\mu$ on $G$ there is a unique real number $\chi(G,\mu)$ such that
@@
Proposition 7.1 ([3, Propositions 1.2 and 3.6], [6, §3.6]).
Let $G$ be a t.d.l.c. group acting on a tree $T$ with compact open vertex stabilisers. Let $\Gamma$ be the quotient graph of $(G,T)$, and denote by $\omega$ its standard edge weight. Then $G$ is unimodular if, and only if, for every closed path $(a_{1},\ldots,a_{n})$ in $\Gamma$ we have
@@
Let $\mathfrak{p}=(a_{1},\ldots,a_{m})$ and $\mathfrak{q}=(b_{1},\ldots,b_{n})$ be reduced paths in $\Gamma$ with $o(a_{1})=o(b_{1})$ and $t(a_{m})=t(b_{n})$. Hence $N_{\mathrm{vert}}(\mathfrak{p})=\prod_{i=1}^{m}\omega(a_{i})$, $N_{\mathrm{edg}}(\mathfrak{p})=1$ if $m=1$ and $N_{\mathrm{edg}}(\mathfrak{p})=\prod_{i=2}^{m}\omega(a_{i})$ if $m\geq 2$. Similar observations hold for $\overline{\mathfrak{p}}$, $\mathfrak{q}$ and $\overline{\mathfrak{q}}$. By Proposition 7.1, we deduce that
@@
Theorem 7.2 ([8, Theorem 5.6]).
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ with compact open vertex stabilisers and finite quotient graph. Let $\mathcal{V}\subseteq VT$ and $\mathcal{E}^{+}\subseteq ET$ be sets of representatives for the $G$-orbits on $VT$ and on a fixed orientation $ET^{+}$ in $T$. Then, for every Haar measure $\mu$ on $G$,
@@
From the hypotheses of Theorem 7.2 one deduces that $T$ is locally finite. In particular, in this case the quantity $\chi(G,\mu)$ coincides with the Euler–Poincaré characteristic of $G$ with respect to $\mu$ as defined in [16, Definition 4.8].
@@
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ with compact open vertex stabilisers, finite quotient graph $\Gamma$, and such that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Let $\omega$ be the standard edge weight on $\Gamma$, and let $N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}$, $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ be as in Definition 4.4. Let $c\in V\Gamma$ and $\Lambda\subseteq\Gamma$ be a maximal subtree, and consider an orientation $E\Lambda^{+}$ in $\Lambda$ such that the restricted origin map $o\colon E\Lambda^{+}\to V\Lambda\setminus\{c\}$ is a bijection. Let also $E\Gamma^{+}$ be an arbitrary orientation in $\Gamma$ such that $E\Gamma^{+}\cap\Lambda=E\Lambda^{+}$.
@@
(7.3) $\chi(G,\mu_{G_{v}})=1+\sum_{a\in E\Gamma^{+}\cap E\Lambda}(1-\omega(a))\frac{N_{\mathrm{vert}}(\mathfrak{p}_{c,o(a)})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}_{c,o(a)}})}-\sum_{b\in E\Gamma^{+}\setminus E\Lambda}\frac{N_{\mathrm{vert}}(\mathfrak{q}_{c,b})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}_{c,b}})},$
for arbitrary reduced paths $\mathfrak{p}_{c,o(a)}\in\mathcal{P}^{\mathrm{lift}}_{\Gamma,v}(c\to o(a))$ and $\mathfrak{q}_{c,b}\in\mathcal{P}^{\mathrm{lift}}_{\Gamma,v}(c\to b)$, for all $a\in E\Gamma^{+}\cap E\Lambda$ and $b\in E\Gamma^{+}\setminus E\Lambda$.
@@
    Let $\Gamma$ be a finite tree. For every $c\in V\Gamma$, there is an orientation $E\Gamma^{+}$ for which the origin map restricts to a bijection $o\colon E\Gamma^{+}\to V\Gamma\setminus\{c\}$.
    Indeed, let $\mathcal{E}^{+}$ be an arbitrary orientation in $\Gamma$ and set
    $E\Gamma^{+}:=\{a\mid a\in\mathcal{E}^{+},\,o(a)\neq c\}\sqcup\{\bar{a}\mid a\in\mathcal{E}^{+},\,o(a)=c\}.$
    Then $E\Gamma^{+}$ is an orientation. Moreover, since $o(a)\neq t(a)$ for every $a\in E\Gamma$, the origin map in $\Gamma$ restricts to a map $o\colon E\Gamma^{+}\to V\Gamma\setminus\{c\}$. By [20, §I.2, Proposition 12] we have $|E\Gamma^{+}|=|V\Gamma|-1$ and then $o\colon E\Gamma^{+}\to V\Gamma\setminus\{c\}$ is bijective.
@@
    By Remark 7.1.1, the right-hand side of (7.3) does not depend on the choice of specific reduced paths $\mathfrak{p}_{c,o(a)}$ and $\mathfrak{q}_{c,b}$ from $c$ to $o(a)$ and from $c$ to $b$, respectively.
@@
Let $e\in\mathcal{E}^{+}$ and set $\pi(e)=a$. For $t\in\{o(e),e\}$, consider the geodesic $[v,t]=(e_{1},\dots,e_{n})$ in $T$ lifting $\mathfrak{p}_{c,o(a)}$ if $t=o(e)$, and lifting $\mathfrak{q}_{c,a}$ if $t=e$. Then,
@@
For the latter equality in (7.4), see (4.5) and (4.6). Note that $|G_{v}\cdot e_{1}|=\omega(\pi(e_{1}))$. Moreover, $|G_{t}\cdot\bar{e}_{n}|=1$ if $t=e$ (because $e_{n}=t$) and $|G_{t}\cdot\bar{e}_{n}|=\omega(\pi(\bar{e}_{n}))$ if $t=o(e)$ (because $o(\bar{e}_{n})=t$). For $11\leq k\leq n-$ 1 we claim that
@@
(7.7) $\frac{1}{\mu_{G_{v}}(G_{t})}=\frac{\omega(\pi(e_{1}))}{|G_{t}\cdot\bar{e}_{n}|}\prod_{i=1}^{n-1}\frac{\omega(\pi(e_{k+1}))}{\omega(\pi(\bar{e}_{k}))},$
where $|G_{t}\cdot\bar{e}_{n}|$ equals $1$ if $t=e$, and it equals $\omega(\pi(\bar{e}_{n}))$ if $t=o(e)$. By design, $\pi([v,t])=:(a_{1},\ldots,a_{n})$ is a reduced path in $\Gamma$. Then
@@
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ with compact open vertex-stabilisers and finite quotient graph $\Gamma$. Denote by $\omega$ the standard edge weight, and assume that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. For every $t\in T$ with $G\cdot t=u$, from Proposition 7.3 we have
@@
    Let $\Gamma$ be a $1$-segment with $E\Gamma=\{a,\bar{a}\}$. Since $\Gamma$ is a tree, $(\Gamma,\omega)$ is unimodular for every $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (cf. Proposition 7.1). Moreover, $\Gamma$ is its only maximal subtree. Consider an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Set $c=o(a)$, $E\Gamma^{+}=\{\bar{a}\}$ and let $\mathfrak{p}_{c,o(\bar{a})}$ be the $1$-edge path $a$. Then Proposition 7.3 implies that
@@
    Let $\Gamma$ be a $n$-bouquet of loops based on the vertex $c$. Note that the $1$-point subgraph is the only maximal subtree of $\Gamma$. By Proposition 7.1, for an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ the pair $(\Gamma,\omega)$ is unimodular if, and only if, $\omega(a)=\omega(\bar{a})$ for every $a\in E\Gamma$. Provided $(\Gamma,\omega)$ is unimodular and $E\Gamma=\{a_{i},\bar{a}_{i}\mid 1\leq i\leq n\}$, from Proposition 7.3 we deduce that
@@
First, (7.12) follows from Remark 7.3.1. By Remark 7.1.1, the ratios in (7.13) and (7.14) do not depend on the choices of $\mathfrak{p}$ and $\mathfrak{q}$, respectively. Moreover, if we prove (7.13) and (7.14) for $\ell(\mathfrak{p})=1$ and $\ell(\mathfrak{q})=2$, the general statements follow iteratively. It remains to observe what follows. First, for every $a\in E\Gamma$ we have
@@
$\begin{split}\chi(\Gamma,c)&=1+\sum_{i=1}^{2}\Bigg{(}\sum_{a\in E\Gamma_{i}^{+}\cap E\Lambda_{i}}(1-\omega(a))\frac{N_{\mathrm{vert}}(\mathfrak{p}_{c,o(a)})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}_{c,o(a)}})}-\sum_{a\in E\Gamma_{i}^{+}\setminus E\Lambda_{i}}\frac{N_{\mathrm{vert}}(\mathfrak{q}_{c,a})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}_{c,a}})}\Bigg{)}\\ &=\chi(\Gamma_{1},c)+\chi(\Gamma_{2},c)-1.\end{split}$
7.2. The evaluation at $s=-1$ and the Euler–Poincaré characteristic
@@
Let $(\Gamma,\omega)$ be a unimodular edge-weighted graph satisfying Setting [$\Gamma$] and such that $\Gamma$ has no cycles of length $\geq 2$. Then, for every $a\in E\Gamma$,
@@
Hence (7.17) follows from (7.23) and (7.24). Finally, assume that both $|o^{-1}(c)|\geq 2$ and $|o^{-1}(d)|\geq 2$. Denote by $\Xi_{1}$ and $\Xi_{2}$ be the connected components of $\Gamma\setminus\{a,\bar{a}\}$ containing $c$ and $d$, respectively. There are exactly two connected components because $\Gamma$ has no cycles of length $\geq 2$. Since $|o^{-1}(c)|\geq 2$ and $|o^{-1}(d)|\geq 2$, both $E\Xi_{1}$ and $E\Xi_{2}$ are non-empty. Moreover, $E\Gamma=E\Xi_{1}\sqcup\{a,\bar{a}\}\sqcup E\Xi_{2}$. For $i\in\{1,2\}$, let $\Gamma_{i}$ be the smallest subgraph of $\Gamma$ containing $\Lambda_{i}\cup\{a,\bar{a}\}$, and note that $k(\Gamma_{i})<k(\Gamma)$. Let also $\Gamma_{a}$ be the $1$-segment subgraph with edge set $\{a,\bar{a}\}$, and observe that $\Gamma_{1}\cap\Gamma_{2}=\Gamma_{a}$. Moreover, if $\Lambda_{1}=\Xi_{1}$ and $\Lambda_{2}=\Gamma_{2}$, we have $\Gamma=\Lambda_{1}\cup\Lambda_{2}$, $\Lambda_{1}\cap\Lambda_{2}=\{c\}$ and $\Lambda_{i}\subseteq\Gamma_{i}$ for every $i\in\{1,2\}$. Hence, Corollary 6.5 and Corollary 6.8 imply
@@
Let $(\Gamma,\omega)$ be a unimodular edge-weighted graph satisfying Setting [$\Gamma$] and such that $\Gamma$ has no cycles of length $\geq 2$. If $\chi(\Gamma,u)=\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for some $u\in\Gamma$, then $\chi(\Gamma,u)=\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for every $u\in\Gamma$.
@@
for some vertex $c\in V\Gamma$. We prove it by induction on $|E\Gamma|/2=:k(\Gamma)\geq 1$. Let first $k(\Gamma)=1$, i.e., $E\Gamma=\{a,\bar{a}\}$. If $o(a)\neq t(a)$, Example 5.14 and Example 7.5(i) yield
@@
Let now $k(\Gamma)\geq 2$ and assume that the statement holds for all graphs $\Gamma^{\prime}$ with $k(\Gamma^{\prime})<k(\Gamma)$. Without loss of generality, we may take $c\in V\Gamma$ such that $|o^{-1}(c)|\geq 2$. Note that this vertex exists because $k(\Gamma)\geq 2$. By Lemma 6.6, there are proper connected subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ such that $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and $\Lambda_{1}\cap\Lambda_{2}=\{c\}$. Then Proposition 6.4 and Lemma 7.7 yield the claim. ∎
@@
7.3. The behaviour at $s=-1$ and the Ihara zeta function of a weighted graph
In [9, §3], a generalisation of the classical Ihara zeta function has been defined for every finite graph $\Gamma$ with a transition weight. Although it is not necessary here, we mention that the finiteness hypothesis on $\Gamma$ can be relaxed. According to [9, Definition 3.3], a transition weight on a finite graph $\Gamma$ is a map $W\colon E\Gamma\times E\Gamma\longrightarrow\mathbb{R}_{\geq 0}$ such that, whenever $W(a,b)\neq 0$, then $t(a)=o(b)$.
According to the definition of graph in [9] (cf. [9, Definition 3.1]), every edge is supposed to be uniquely determined by its endpoints. However, for the results involved below, this hypothesis has no influence and thus we do not assume it.
The Ihara zeta function $Z_{(\Gamma,W)}(x)$ of $(\Gamma,W)$ has been defined in [9, Definition 3.8] as a suitable infinite product of meromorphic functions on $\mathbb{C}$ converging for all $x\in\mathbb{C}$ with $|x|\ll 1$. Here we only need the following characterisation of the reciprocal of $Z_{(\Gamma,W)}(x)$, cf. [9, Theorem 3.11]. Namely,

### v2

@@ 6. The reciprocal of $\mathcal{Z}_{\Gamma,u\to u}(s)$
7. The behaviour at s=1s=-1
The main goal of this section is to prove that the Euler–Poincaré identity (1.3) holds for a unimodular t.d.l.c. group $G$ having an action on a tree as prescribed by Corollary F. To achieve this, we first give a formula of the relevant Euler–Poincaré characteristic in terms of local data of the action (cf. Proposition 7.3) and then use the splitting formulae of Section 6 to prove Theorem E and Corollary F. Finally, given an edge-weighted graph $(\Gamma,\omega)$, we exploit the machinery introduced in this paper to relate the behaviour of $\mathcal{Z}_{\Gamma,u\to u}(-1)$ with the behaviour at $1$ of a suitable weighted Ihara zeta function associated to $\Gamma$ (cf. Section 7.3).
@@ 7.1. The Euler–Poincaré characteristic
According to [10, §5], every unimodular t.d.l.c. group $G$ of type $\mathrm{FP}$ (with respect to the category of discrete left $\mathbb{Q}[G]$-modules) admits an Euler–Poincaré characteristic $\widetilde{\chi}_{G}$. For every compact open subgroup $K\leq G$, this invariant is a determined rational multiple of the Haar measure $\mu_{K}$ on $G$ normalised with respect to $K$, written $\widetilde{\chi}_{G}=\chi(G,\mu_{K})\cdot\mu_{K}$. By the uniqueness of the Haar measure on $G$ up to a positive real rescaling, for every Haar measure $\mu$ on $G$ there is a unique real number $\chi(G,\mu)$ such that
@@
Proposition 7.1 ([3, Propositions 1.2 and 3.6], [8, §3.6]).
Let $G$ be a t.d.l.c. group acting on a tree $T$ with compact open vertex stabilisers and without edge inversions. Let $\Gamma$ be the quotient graph of $(G,T)$ and denote by $\omega$ its standard edge weight. Then $G$ is unimodular if and only if, for every closed path $(a_{1},\ldots,a_{n})$ in $\Gamma$,
@@
Let $\mathfrak{p}=(a_{1},\ldots,a_{m})$ and $\mathfrak{q}=(b_{1},\ldots,b_{n})$ be reduced paths in $\Gamma$ with $o(a_{1})=o(b_{1})$ and $t(a_{m})=t(b_{n})$. Hence $N_{\mathrm{vert}}(\mathfrak{p})=\prod_{i=1}^{m}\omega(a_{i})$. Moreover, $N_{\mathrm{edg}}(\mathfrak{p})=1$ if $m=1$ and $N_{\mathrm{edg}}(\mathfrak{p})=\prod_{i=2}^{m}\omega(a_{i})$ if $m\geq 2$. Similar observations hold for $\overline{\mathfrak{p}}$, $\mathfrak{q}$ and $\overline{\mathfrak{q}}$. By Proposition 7.1, we deduce that
@@
Theorem 7.2 ([10, Theorem 5.6]).
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ without edge inversions, with compact open vertex stabilisers and with finite quotient graph. Let $\mathcal{V}\subseteq VT$ and $\mathcal{E}^{+}\subseteq ET$ be sets of representatives for the $G$-orbits on $VT$ and on a fixed orientation $ET^{+}$ in $T$, respectively. Then, for every Haar measure $\mu$ on $G$,
@@
The hypotheses of Theorem 7.2 imply that $T$ is locally finite. In particular, in this case the quantity $\chi(G,\mu)$ coincides with the Euler–Poincaré characteristic of $G$ with respect to $\mu$ as defined in [20, Definition 4.8].
@@
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$ without edge inversions, with compact open vertex stabilisers, with finite quotient graph $\Gamma$, and such that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. Let $\omega$ be the standard edge weight on $\Gamma$, and let $N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}$, $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ be as in Definition 4.4. Let $c\in V\Gamma$ and $\Lambda\subseteq\Gamma$ be a maximal subtree, and consider an orientation $E\Lambda^{+}$ in $\Lambda$ such that the restricted origin map $o\colon E\Lambda^{+}\to V\Lambda\setminus\{c\}$ is a bijection. Let also $E\Gamma^{+}$ be an arbitrary orientation in $\Gamma$ such that $E\Gamma^{+}\cap\Lambda=E\Lambda^{+}$.
@@
(7.3) $\chi(G,\mu_{G_{v}})=1+\sum_{a\in E\Lambda^{+}}(1-\omega(a))\frac{N_{\mathrm{vert}}(\mathfrak{p}_{c,o(a)})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}_{c,o(a)}})}-\sum_{b\in E\Gamma^{+}\setminus E\Lambda}\frac{N_{\mathrm{vert}}(\mathfrak{q}_{c,b})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}_{c,b}})},$
for arbitrary reduced paths $\mathfrak{p}_{c,o(a)}\in\mathcal{P}_{\Gamma,v}(c\to o(a))$ and $\mathfrak{q}_{c,b}\in\mathcal{P}_{\Gamma,v}(c\to b)$, for all $a\in E\Lambda^{+}$ and $b\in E\Gamma^{+}\setminus E\Lambda$.
@@
    Let $\Lambda$ be a finite tree. For every $c\in V\Lambda$, there is an orientation $E\Lambda^{+}$ for which the origin map restricts to a bijection $o\colon E\Lambda^{+}\to V\Lambda\setminus\{c\}$.
    Indeed, let $\mathcal{E}^{+}$ be an arbitrary orientation in $\Lambda$ and set
    $E\Lambda^{+}:=\{a\mid a\in\mathcal{E}^{+},\,o(a)\neq c\}\sqcup\{\bar{a}\mid a\in\mathcal{E}^{+},\,o(a)=c\}.$
    Then $E\Lambda^{+}$ is an orientation. Moreover, since $o(a)\neq t(a)$ for every $a\in E\Lambda$, the origin map in $\Lambda$ restricts to a map $o\colon E\Lambda^{+}\to V\Lambda\setminus\{c\}$. By [27, §I.2, Proposition 12] we have $|E\Lambda^{+}|=|V\Lambda|-1$ and then $o\colon E\Lambda^{+}\to V\Lambda\setminus\{c\}$ is bijective.
@@
    By Remark 7.1.1, the right-hand side of (7.3) does not depend on the choice of specific reduced paths $\mathfrak{p}_{c,o(a)}$ and $\mathfrak{q}_{c,b}$ from $c$ to $o(a)$ and from $c$ to $b$, respectively. Being $\Gamma$ connected, note that there is always a reduced path connecting any two vertices of $\Gamma$. Reduced paths in $\Gamma$ are always liftable to geodesics in $T$ (even if $T$ has leaves).
@@
Let $e\in\mathcal{E}^{+}$ and $\pi(e)=a$. For $t\in\{o(e),e\}$, consider the geodesic $v,t]=(e_{1},\dots,e_{n})$ in $T$ lifting $\mathfrak{p}_{c,o(a)}$ if $t=o(e)$, and lifting $\mathfrak{q}_{c,a}$ if $t=e$. Then,
@@
For the latter equality in (7.4), see (4.5) and (4.6). Note that $|G_{v}\cdot e_{1}|=\omega(\pi(e_{1}))$. Moreover, $|G_{t}\cdot\bar{e}_{n}|=1$ if $t=e$ (because $e_{n}=t$) and $|G_{t}\cdot\bar{e}_{n}|=\omega(\pi(\bar{e}_{n}))$ if $t=o(e)$ (because $o(\bar{e}_{n})=t$). For $1\leq k\leq n-1$ we claim that
@@
(7.7) $\frac{1}{\mu_{G_{v}}(G_{t})}=\frac{\omega(\pi(e_{1}))}{|G_{t}\cdot\bar{e}_{n}|}\prod_{k=1}^{n-1}\frac{\omega(\pi(e_{k+1}))}{\omega(\pi(\bar{e}_{k}))},$
where $|G_{t}\cdot\bar{e}_{n}|$ equals 11 if $t=e$, and it equals $\omega(\pi(\bar{e}_{n}))$ if $t=o(e)$. By design, $\pi([v,t])=:(a_{1},\ldots,a_{n})$ is a reduced path in $\Gamma$. Then
@@
Let $G$ be a unimodular t.d.l.c. group acting on a tree $T$, without edge inversions, with compact open vertex-stabilisers and with finite quotient graph $\Gamma$. Denote by $\omega$ the standard edge weight, and assume that $(G,T)$ is weakly locally $\infty$-transitive or (P)-closed. For every $t\in T$ with $G\cdot t=u$, from Proposition 7.3 we have
@@
    Let $\Gamma$ be a $1$-segment with $E\Gamma=\{a,\bar{a}\}$. Since $\Gamma$ is a tree, $(\Gamma,\omega)$ is unimodular for every $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (cf. Proposition 7.1). Moreover, $\Gamma$ is its only maximal subtree. Consider an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Set $c=o(a)$, $E\Gamma^{+}=\{\bar{a}\}$ and let $\mathfrak{p}_{c,o(\bar{a})}$ be the 11-edge path $a$. Then Proposition 7.3 implies that
@@
    Let $\Gamma$ be a $n$-bouquet of loops based on the vertex $c$. Note that the 11-point subgraph is the only maximal subtree of $\Gamma$. By Proposition 7.1, for an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ the pair $(\Gamma,\omega)$ is unimodular if, and only if, $\omega(a)=\omega(\bar{a})$ for every $a\in E\Gamma$. Provided $(\Gamma,\omega)$ is unimodular and $E\Gamma=\{a_{i},\bar{a}_{i}\mid 1\leq i\leq n\}$, from Proposition 7.3 we deduce that
@@
First, (7.12) follows from Remark 7.3.1(iii). By Remark 7.1.1, the ratios in (7.13) and (7.14) do not depend on the choices of $\mathfrak{p}$ and $\mathfrak{q}$, respectively. Moreover, if we prove (7.13) and (7.14) for $\ell(\mathfrak{p})=1$ and $\ell(\mathfrak{q})=2$, the general statements follow iteratively. It remains to observe what follows. First, for every $a\in E\Gamma$ we have
@@
$\begin{split}\chi(\Gamma,c)&=1+\sum_{i=1}^{2}\Bigg(\sum_{a\in E\Gamma_{i}^{+}\cap E\Lambda_{i}}\hskip-5.69046pt(1-\omega(a))\frac{N_{\mathrm{vert}}(\mathfrak{p}_{c,o(a)})}{N_{\mathrm{vert}}(\overline{\mathfrak{p}_{c,o(a)}})}-\hskip-11.38092pt\sum_{a\in E\Gamma_{i}^{+}\setminus E\Lambda_{i}}\hskip-5.69046pt\frac{N_{\mathrm{vert}}(\mathfrak{q}_{c,a})}{N_{\mathrm{edg}}(\overline{\mathfrak{q}_{c,a}})}\Bigg)\\ &=\chi(\Gamma_{1},c)+\chi(\Gamma_{2},c)-1.\end{split}$
7.2. The evaluation at s=1s=-1 and the Euler–Poincaré characteristic
@@
Let $(\Gamma,\omega)$ be a unimodular edge-weighted graph satisfying Setting $\Gamma$] and such that $\Gamma$ has no cycles of length $\geq 2$. Then, for every $a\in E\Gamma$,
@@
Hence (7.17) follows from (7.23) and (7.24). Finally, assume that both $|o^{-1}(c)|\geq 2$ and $|o^{-1}(d)|\geq 2$. Denote by $\Xi_{1}$ and $\Xi_{2}$ the connected components of $\Gamma\setminus\{a,\bar{a}\}$ containing $c$ and $d$, respectively. There are exactly two connected components because $\Gamma$ has no cycles of length $\geq 2$. Since $|o^{-1}(c)|\geq 2$ and $|o^{-1}(d)|\geq 2$, both $E\Xi_{1}$ and $E\Xi_{2}$ are non-empty. Moreover, $E\Gamma=E\Xi_{1}\sqcup\{a,\bar{a}\}\sqcup E\Xi_{2}$. For $i\in\{1,2\}$, let $\Gamma_{i}$ be the smallest subgraph of $\Gamma$ containing $\Xi_{i}\cup\{a,\bar{a}\}$, and note that $k(\Gamma_{i})<k(\Gamma)$. Let also $\Gamma_{a}$ be the $1$-segment subgraph with edge set $\{a,\bar{a}\}$, and observe that $\Gamma_{1}\cap\Gamma_{2}=\Gamma_{a}$. Moreover, if $\Lambda_{1}:=\Xi_{1}$ and $\Lambda_{2}:=\Gamma_{2}$, we have $\Gamma=\Lambda_{1}\cup\Lambda_{2}$, $\Lambda_{1}\cap\Lambda_{2}=\{c\}$ and $\Lambda_{i}\subseteq\Gamma_{i}$ for every $i\in\{1,2\}$. Hence, Corollary 6.5 and Corollary 6.8 imply
@@
Let $(\Gamma,\omega)$ be a unimodular edge-weighted graph satisfying Setting $\Gamma$] and such that $\Gamma$ has no cycles of length $\geq 2$. If $\chi(\Gamma,u)=\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for some $u\in\Gamma$, then $\chi(\Gamma,u)=\mathcal{Z}_{\Gamma,u\to u}(-1)^{-1}$ for every $u\in\Gamma$.
@@
for some vertex $c\in V\Gamma$. We prove (7.25) by induction on $|E\Gamma|/2=:k(\Gamma)\geq 1$. Let first $k(\Gamma)=1$, i.e., $E\Gamma=\{a,\bar{a}\}$. If $o(a)\neq t(a)$, Example 5.16 and Example 7.5(i) yield
@@
Let now $k(\Gamma)\geq 2$ and assume that the statement holds for all graphs $\Gamma^{\prime}$ with $k(\Gamma^{\prime})<k(\Gamma)$. Without loss of generality (cf. Corollary 7.9), we may take $c\in V\Gamma$ such that $|o^{-1}(c)|\geq 2$. Note that this vertex exists because $k(\Gamma)\geq 2$. By Lemma 6.6, there are proper connected subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=\{c\}$. Then Proposition 6.4 and Lemma 7.7 yield the claim. ∎
@@
7.3. The behaviour at s=1s=-1 and the Ihara zeta function of a weighted graph
In [11, §3], a generalisation of the classical Ihara zeta function has been defined for every finite graph $\Gamma$ with a transition weight. Although it is not necessary here, we mention that the finiteness hypothesis on $\Gamma$ can be relaxed. According to [11, Definition 3.3], a transition weight on a finite graph $\Gamma$ is a map $W\colon E\Gamma\times E\Gamma\longrightarrow\mathbb{R}_{\geq 0}$ such that, whenever $W(a,b)\neq 0$, then $t(a)=o(b)$.
According to the definition of graph in [11] (cf. [11, Definition 3.1]), every edge is supposed to be uniquely determined by its endpoints. However, for the results involved below, this hypothesis has no influence and thus we do not assume it.
The Ihara zeta function $Z_{(\Gamma,W)}(x)$ of $(\Gamma,W)$ has been defined in [11, Definition 3.8] as a suitable infinite product of meromorphic functions on $\mathbb{C}$ converging for all $x\in\mathbb{C}$ with $|x|\ll 1$. Here we only need the following characterisation of the reciprocal of $Z_{(\Gamma,W)}(x)$, cf. [11, Theorem 3.11]. Namely,